OpenStudy (anonymous):

find the second derivation : y=(x^3 - 7x + 1)^4

OpenStudy (espex):

Where are you stuck?

OpenStudy (anonymous):

first order derivative is 4(x^3-7x+1)^3 (3x^2-7)

OpenStudy (anonymous):

now using multipplication rule again differentiate it wrt x !

OpenStudy (nincompoop):

there are two ways to solve this. first, you can expand then do a normal power rule on each term the other method is to use chain rule

OpenStudy (nincompoop):

once you've done one of the two ways, you can just go ahead and obtain the second derivative

OpenStudy (anonymous):

\[\huge y=(x^3 - 7x + 1)^4\] \[\huge y'=4(x^3 - 7x + 1)^3 (3x^2 -7)\] \[ y''=4(x^3 - 7x + 1)^3 (6x)+4(3x^2 -7)3(3x^2 - 7 + 1)^2 (6x)\] Now simplify it

OpenStudy (anonymous):

then the second derivative is = 3*4(x^3-7x+1)^2(3x^2-7)(3x^2-7)+4(x^3-7x+1)(6x)

OpenStudy (anonymous):

thanks a lot for all :)

OpenStudy (anonymous):

y′′=4(x3−7x+1)3(6x)+4(3x2−7)3(3x2−7+1)2(6x) I think this it's wrong :(

OpenStudy (anonymous):

ya it is ! why multiplying 3 in the first term ? I think the correct ans =3*4(x^3-7x+1)^2(3x^2-7)(3x^2-7)+4(x^3-7x+1)(6x)

OpenStudy (anonymous):

12(x^3-7x+1)^2(3x^2-7)+4(x^3-7x+1)^3(6x) this is correct ?

OpenStudy (anonymous):

12(x^3-7x+1)^2(3x^2-7)(3x^2-7)+4(x^3-7x+1)^3(6x) sorry I mean this .. ?

OpenStudy (anonymous):

ohh yea

OpenStudy (anonymous):

thanks @moli1993 :)

OpenStudy (anonymous):

np, :)